22-1 Calculate the electrode potentials of the following half-cells. Ag (0.0436M)|Ag a. b. Fe3 (5.34x10-4M), Fe2 (0.090M)|Pt C.
Answer #1

a. Ag^+ (aq) + e^- ------------------> Ag(s)     E0 = 0.80v

n = 1

Ecell   =    E0cell -0.0592/n logQ

            = 0.80 -0.0592/1log1/[Ag^+]

           = 0.80-0.0592log1/0.0436

         = 0.80-0.0592*.13606   = 0.79v

b. Fe^3+ (aq) + e^- --------> Fe^2+ (aq)       E0 = 0.77v

n = 1

Ecell   =    E0cell -0.0592/n logQ

            = 0.77 -0.0592/1log[Fe^2+]/[Fe^3+]

           = 0.77-0.0592log0.09/0.000534

            = 0.77-0.0592*2.2267   = 0.64v

c. AgBr(s) + e^----------------------> Ag(s) + Br^-     E0 = 0.0713v

n =1

Ecell   =    E0cell -0.0592/n logQ

            = 0.0713 -0.0592/1log[Br^-]

           = 0.0713-0.0592log0.037

        = 0.0713-0.0592*-1.4317    = 0.156v >>>>answer